OpenStudy (beccaboo333):

Solve the following system using the substitution method. x + 6y = -5 2x - 9y = -17

OpenStudy (beccaboo333):

it's a knowledge thing in a lesson. the before the actual assignment. i'm not good with math. i need help. can someone explain?

hero (hero):

No need to post answer choices for this one.

OpenStudy (beccaboo333):

alright hero, sorry

OpenStudy (anonymous):

alrigh becca this is the same as the other one

hero (hero):

So to solve by substitution, first isolate one of the variables. Which variable do you think would be most appropriate to isolate @beccaboo333

OpenStudy (beccaboo333):

the x ones?

OpenStudy (anonymous):

x+6y=-5 -6y -6y x=-6y-5 2(-6y-5)-9y=-17

OpenStudy (beccaboo333):

what what?

OpenStudy (anonymous):

look the 6y go down in transfort into a - just to moved to the other side

OpenStudy (beccaboo333):

what? im sorry. i literally do not understand this..

OpenStudy (anonymous):

u basically want to get a variable alone on one side and then substitute whatever its equal to into the other equation to get your answer is what hes trying to show

OpenStudy (anonymous):

so he took x=-6y-5 and substitute it into the other equation wherever their was a x present so you could then solve because their would only be 1 variable

OpenStudy (beccaboo333):

uhm, still.. what?

OpenStudy (anonymous):

so he took the equation x+6y=-5 he added 6y to both sides to get the x alone on the left x=-6y-5 so its sayin x is equal to -6y-5 you then take the other equation 2x-9y= -17 and wherever you see an x u replace it with your equatrion so then its 2(-6y-5)-9y=-17 by doing so u can then solve for y you distribute the 2 giving you -12y-10 -9y=-17 combine like terms u get -21y-10=-17 then add 10 to both sides giving u -21y=-7 divide by 21 so u get y=7/21 that simplifies to 1/3 finally u plug y into either original equation(whichever is easier) to solve for x so u get x+6(1/3)=-5 u multiply 6/1*1/3 so u get x+2/1=-5 2/1= 2 so its x+2=-5 subtract 2 on both sides giving u x= -7 your final answer is y=1/3 x= -7

OpenStudy (anonymous):

explained every step hopefully u get it now

OpenStudy (beccaboo333):

omg.. i love you so much. thank you!!

OpenStudy (anonymous):

So U understand it now? cuz if not maybe i can explain it some other way

OpenStudy (beccaboo333):

it's \[-7 1/3 \]

OpenStudy (beccaboo333):

there's supposed to be a space

OpenStudy (beccaboo333):

but yes i do

OpenStudy (anonymous):

oops srry but the steps should be the same

OpenStudy (beccaboo333):

so -7 and 1/3?

OpenStudy (anonymous):

it should be not 2 sure what u ment by -71/3

OpenStudy (beccaboo333):

nothing. thanks!

OpenStudy (anonymous):

okay no problem glad 2 help

OpenStudy (anonymous):

is A BECCA

OpenStudy (anonymous):

wait noo it was B SOORRYY

OpenStudy (anonymous):

@beccaboo333